I am studying for a final and I'm really having a difficult time with functions. I don't know why I don't understand them, but I don't. Please help with the following:
Given f(x)= 1/sqrt of x and g(x) = 1-x^2 evaluate the following:
a) g(2a+1)
b) g(f(x))
Thank you for your help.

Answer by stanbon(48558)

Given f(x)= 1/sqrt of x and g(x) = 1-x^2 evaluate the following:
a) g(2a+1)
Substitute "2a+1" for x to get:
g(2a+1) = 1 - (2a+1)^2
=1 -[4a^2+4a+1]
= -4a^2-4a
------------
b) g(f(x))
=g[(1/sqrt(x)]
= 1 - (1/sqrt(x))^2
= 1 - (1/x)
-------------
Cheers,
Stan H.